OpenStudy (anonymous):

using the replacement set, find the solutions for the equation. y=0.5x+10 replacement set (-2, 9.5) , (0,10) , (3,11.5) ,(4, 13)

OpenStudy (anonymous):

I think you should have made 8 questions out of this. Here is the first one: The table below represents the relationship between the time Haley is typing on the computer and the number of pages she has typed. Which is the dependent variable: time or pages? Time (min) 10 20 30 40 50 Number of Pages 1 2 3 4 5 A) Time B) Pages usually the independent variable is listed first followed by the dependent variable. that would make the Number of Pages the dependent variable.

hero (hero):

Hint: (0,10) is one of the solutions y = 0.5x + 10 10 = 0.5(0) + 10 10 = 10

hero (hero):

There is another solution to this by the way